OpenStudy (zenmo):

Does anyone know how to use a ti-83 graphing calculator for an arc length approximation?

OpenStudy (zenmo):

\[\int\limits_{0}^{5}\sqrt{1+\frac{ 9x^4 }{ (6+x^3)^2 }}dx\]

OpenStudy (zenmo):

I need to use a calculator to approximate this, but don't know how so.